# Tensor-Var: Variational Data Assimilation in Tensor Product Feature Space

### v2

@@ 4.2 Global NWP
Next, we consider a global NWP problem. The European Centre for Medium-Range Weather Forecasts (ECMWF) Atmospheric Reanalysis (ERA5) dataset (Hersbach et al., 2020) provides the best estimate of the dynamics of the atmosphere covering the period from 1940 - present. For a proof of concept, the 500hPa geopotential, 850 hPa temperature, 700 hPa humidity, and 850 hPa wind speed (meridonal and zonal directions) at 64$\times$32 resolution is considered here. The data is sourced from the WeatherBench2 repository (Rasp et al., 2024). Observations are sampled randomly from the grid with a 15% spatial coverage and with additive noise (0.01 times the standard deviation of the state variable) (see Figure 2). Latent dimensions of the two baseline methods are set to match the feature dimension $d_{s}$ in our approach. The evaluation metric is the area-weighting RMSE over grid points (see more details in Appendix C.3). We trained all the models on ERA5 data from 1979-01-01 to 2016-01-01 and tested on data post-2018, with a qualitative evaluation shown for 2018-01-01 00:00 in Figure 2.
Figure 3 (left 1-5) shows the distributions of NRMSE (%) across different atmospheric variables (z500, t850, q700, u850, v850) for Latent 3D-Var, Latent 4D-Var, and Tensor-Var, evaluated on a test dataset consisting two years of data from 2018-01-01 00:00 to 2020-01-01 00:00. Our approach consistently achieves the lowest mean and standard deviation of NRMSE for all variables, demonstrating improved performance in both assimilation accuracy and robustness. The latent space of 3D-Var and 4D-Var is learned by autoencoder purely based on the reconstruction loss without considering system dynamics. This weakens the forecasting abilities of the dynamical systems, introducing extra errors in the 4D-Var optimization compared to the 3D-Var. In contrast, our approach jointly learns the feature space representation and linear system dynamics, resulting in more accurate forecasting and improved 4D-Var performances.
The rightmost bar plot in Figure 3 shows evaluation times on an Nvidia RTX-4090 GPU. Tensor-Var outperforms Latent 3D-Var and Latent 4D-Var, demonstrating its better computational efficiency over the data-driven Var-DA approaches. This is because our deep features are used only for mapping data into the feature space rather than being directly involved in the gradient-based optimization via AD. In addition to the assimilation results, we evaluate the forecasting quality of Tensor-Var based on the assimilated state, quantatitive results can be found in Appendix C.3 Figure 10.
Figure 2: Visualization of assimilation results for five variables from ERA5 data at time 2018-01-01 00:00. Each column (from left to right) displays the background state, observations, true state, and errors for Latent-3DVar, Latent-4DVar, and Tensor-Var. The reported error was a weighted absolute error in each pixel; see Appendix C.3 for more details.
Figure 3: Comparison of distribution of NRMSE (%) across different atmospheric variables (z500, t850, q700, u850, v850) for Latent 3D-Var, Latent 4D-Var, and Tensor-Var. The rightmost bar plot shows evaluation times and error bars indicate the standard deviation for evaluation time.
@@ 4.3 Assimilation from Satellite Observations
In the final experiment, in contrast to the randomly selected observation locations in section 4.2, we now consider a more realistic DA problem in global NWP by incorporating the location of satellite tracks. Weather satellites are a critical source of observational data for DA in global NWP; whilst introducing this observation is nontrivial as it incorporates further data fusion of the dynamic observation locations of the satellites and the spatial-temporal sparsity of satellite tracks. The spatial resolution of the underlying grid is even increased to $240\times 121$ in this case.
An example of satellite tracks and observation distribution is shown in Figure 4. We extract satellite track data (latitude and longitude coordinates) from CelesTrak [^2] CelesTrak provides publicly accessible orbital data for a wide range of satellites, including those with meteorological sensors at www.celestrak.com. The data include positional details, and temporal information, allowing for accurate real-time satellite tracking. for the same periods as Section 4.2, matching it with ERA5 data to generate practical observations. These observations include satellite locations within two hours before the assimilation time, sampled at half-hour intervals, with an average coverage of approximately 6%, see Figure 6. Other experimental settings, such as data volume, training/testing periods, variables, and the 4D-Var window, align with Section 4.2.
Figure 4: Example of selected satellite tracks over a one-hour horizon, with observations (black triangles) sampled at half-hour intervals.
In this experiment, we evaluate our proposed method in an online operational scenario, where the DA is continuously applied. To assess the uncertainties, all methods are evaluated 10 times by randomly sampling sequences from the test dataset. We present the mean and standard deviation of NRMSEs in Figure 5 over a time-horizon of 7-days. Tensor-Var consistently achieves the lowest mean and standard deviation, demonstrating its robustness in a large-scale system with practical observations. Figure 6 presents the qualitative assimilation results for the variable z500 with observations; results for other variables are provided in Appendix C.4. Dynamic satellite observations impact assimilation accuracy, with clustered observations near the equator from geosynchronous satellites substantially reducing errors in the corresponding regions. The results demonstrate that Tensor-Var excels in accuracy and robustness, when handling large-scale systems with practical observations, showing its potential for applications in operational DA and forecasting.
Figure 5: Comparison of assimilation NRMSE (%) over a 7-day horizon for five atmospheric variables from Latent 3D-Var, Latent 4D-Var, and Tensor-Var. Each time-step represents a 6-hour interval.
@@ 4.4 Ablation study
To support our empirical results, we conducted ablation studies on the 40- and 80-dimensional Lorenz-96 system to investigate (1) the effect of history length in learning the operator $\hat{\mathcal{C}}_{S|OH}$, and to explore (2) the effect of DFs $\phi_{S}$ with different feature dimensions.
@@
We explored the effect of history length $m$ on learning the inverse observation operator and its impact on state estimation accuracy. According to the theory in (Liu et al., 2022), the history length can be chosen as $m\propto\log(n_{s})$ [^3] Please note that the result omits the class of systems with exponential dependency on the history length.. The ablation study was conducted by scaling $m$ proportionally to $m\approx C\log{n_{s}}$ where the constant $C$ was adjusted. The characteristic dimensions $d_{s},d_{o},d_{h}$ are fixed to be the same as the experiments in Subsection 4.1.
Table 3 shows that incorporating history ($C>0$) significantly improves the accuracy of the state estimation, with the NRMSE decreasing as $C$ increases. However, the improvements become marginal when increasing $C$ beyond a certain point (around $C=4$), suggesting diminishing improvements for longer history. This indicates a trade-off, where increasing the history length beyond a certain threshold yields little additional benefit in state estimation.
Table 2: Comparison of different history lengths, with NRMSE as the metric for state estimation accuracy. $C=0$ indicates that no history is incorporated.
$C=0$ $C=1$ $C=2$ $C=4$ $C=8$
$n_{s}=40$ $11.7\pm 3.4$ $9.3\pm 2.8$ $7.7\pm 1.4$ $7.7\pm 0.9$ $7.5\pm 0.5$
$n_{s}=80$ $10.8\pm 2.6$ $8.2\pm 1.7$ $7.4\pm 0.9$ $7.1\pm 0.9$ $7.0\pm 0.8$
@@
In this experiment, we compared the DFs with different dimensions $d_{s}=20,40,60,80$ and Gaussian kernel feature. To scale up the Gaussian kernel, we applied Nyström approximation and kernel PCA (See Appendix C.5 for details). All three features are Gaussian kernels. The pre-image of the system state was learned using kernel ridge regression on the low-dimensional representations. As shown in Table 3, Gaussian kernel performance was close to the best DFs in $n_{s}=40$ but degraded in $n_{s}=80$. In lower-dimensional problems, pre-determined kernel features were more robust than DFs, but their performance can degrade with increasing system dimensionality and complexity. DFs with $d_{s}=60,120$ consistently performed well, while $d_{s}=20,40$ performed poorly, reflecting the trade-offs among data size, parameter count, and system dimension.
Table 3: Comparison of different features. The metric is NRMSE (%) over the assimilation window
$d_{s}=20$ $d_{s}=40$ $d_{s}=60$ $d_{s}=120$ Gaussian kernel
$n_{s}=40$ $16.7\pm 2.1$ $14.1\pm 1.3$ $\mathbf{8.3}\pm 0.9$ $9.7\pm 0.7$ $8.5\pm\mathbf{0.5}$
$n_{s}=80$ $17.3\pm 2.6$ $16.8\pm 2.2$ $11.4\pm 0.9$ $\mathbf{9.0\pm 0.9}$ $14.3\pm 1.4$
@@ 5 Conclusion
In this paper, we introduced the Tensor-Var method, using kernel conditional mean embedding to transform nonlinear dynamical and observation models in 4D-Var into a linear framework, making the optimization both tractable and computationally efficient. By learning adaptive deep features, Tensor-Var addresses the scalability typically associated with traditional kernel methods. Our inverse observation operator, which incorporates historical observations, improving the accuracy and robustness with incomplete observations. Experiments on two chaotic systems and global weather forecasting show that Tensor-Var outperforms state-of-the-art hybrid ML-DA models in both accuracy and efficiency.
Limitations.
@@
Our method requires access to exact or re-analyzed system states to learn the dynamical and observation models, which may not be feasible in practical applications. A future direction would be to learn these models directly from observations and calibrate the learned dynamics in the feature space using Tensor-Var. In addition, our simplified error covariance matrices in 4D-Var may not fully capture system correlations and uncertainties. Future work could focus on improving the design of error covariance matrices within the Tensor-Var framework to enhance assimilation and performance.
Appendix A Theoretical Analysis

### v3

@@ 4.2 Global NWP
Next, we consider a global NWP problem. The European Centre for Medium-Range Weather Forecasts (ECMWF) Atmospheric Reanalysis (ERA5) dataset (Hersbach et al., 2020) provides the best estimate of the dynamics of the atmosphere covering the period from 1940 to present. The 500hPa geopotential, 850 hPa temperature, 700 hPa humidity, and 850 hPa wind speed (meridional and zonal directions) at 64$\times$32 resolution are considered here. The data is sourced from the WeatherBench2 repository (Rasp et al., 2024). Observations are sampled randomly from the grid with a 15% spatial coverage and with additive noise (0.01 times the standard deviation of the state variable) (see Figure 9). Latent dimensions of the two baseline methods are set to match the feature dimension $d_{s}$ in our approach. The evaluation metric is the area-weighting RMSE over grid points (see more details in Appendix E.3). We trained all models in ERA5 data from 1979-01-01 to 2016-01-01 and tested on data post-2018, with a qualitative evaluation shown for 2018-01-01 00:00 in Figure 9.
Figure 2 (left 1-5) shows the distributions of NRMSE (%) across different atmospheric variables (z500, t850, q700, u850, v850) for Latent 3D-, 4D-Var, and Tensor-Var, evaluated on a test data set consisting of two years of data from 2018-01-01 00:00 to 2020-01-01 00:00. Our approach consistently achieves the lowest mean and standard deviation of NRMSE for all variables, demonstrating improved performance in both assimilation accuracy and robustness. The latent spaces of 3D- and 4D-Var are learned by the autoencoder purely based on the reconstruction loss without considering the system dynamics. This weakens the forecasting abilities of the dynamical systems, introducing extra errors in the 4D-Var optimization compared to the 3D-Var optimization. In contrast, our approach jointly learns feature space representation and linear system dynamics, resulting in more accurate forecasting and improved 4D-Var performances. The rightmost barplot in Figure 2 shows the evaluation times on an Nvidia RTX-4090 GPU. Tensor-Var achieves faster computation than Latent 3D-Var and Latent 4D-Var because its deep features are used only to map data into feature space rather than being directly involved in gradient-based optimization via AD. In addition to the assimilation results, we evaluate the forecast quality of Tensor-Var based on the assimilated state; quantitative results can be found in Appendix E.3 Figure 10.
@@ 4.3 Assimilation from Satellite Observations
In the final experiment, we consider observations from satellite track locations instead of random observation points in Section 4.2, addressing a more realistic DA problem in global NWP. Weather satellites offer crucial observations, but their dynamic positions and spatial-temporal sparsity make data assimilation more challenging. The spatial resolution of the grid increases to $240\times 121$ in this case.
Figure 4: Selected satellite tracks over a one-hour horizon, with observations (black triangles) sampled at half-hour intervals.
An example of satellite tracks and observation distribution is shown in Figure 4. We extract satellite track data (latitude and longitude coordinates) from CelesTrak [^2] CelesTrak provides public orbital data for a wide range of satellites, including those with meteorological sensors at www.celestrak.com. The data include positional details and temporal information, allowing accurate real-time satellite tracking. for the same periods as Section 4.2, matching it with ERA5 data to generate practical observations. These observations include satellite locations within two hours before the assimilation time, sampled at half-hour intervals, with an average coverage of approximately 6%, see Figure 5. Other experimental settings, such as data volume, training/testing periods, variables, and the 4D-Var window, align with Section 4.2.
In this experiment, we evaluate our proposed method in an online operational scenario, where the DA is continuously applied. To assess the uncertainties, all methods are evaluated 10 times by randomly sampling sequences from the test dataset. We present the mean and standard deviation of the NRMSEs in Figure 3 over a 7-day time horizon. Tensor-Var consistently achieves the lowest mean and standard deviation, demonstrating its robustness in a large-scale system with practical observations. Figure 5 presents the qualitative assimilation results for the variable z500 with observations; results for other variables are provided in Appendix E.4. Dynamic satellite observations impact assimilation accuracy, with clustered observations near the equator from geosynchronous satellites substantially reducing errors in the corresponding regions. The results demonstrate that Tensor-Var excels in accuracy and robustness when handling large-scale systems with practical observations, showing its potential for applications in operational DA and forecasting.
@@ 4.4 Ablation study
To support our empirical results, we perform ablation studies on the 40- and 80-dimensional Lorenz-96 system to investigate (1) the effect of history length in learning the operator $\hat{\mathcal{C}}_{S|OH}$, and explore (2) the effect of DFs $\phi_{S}$ with different feature dimensions.
@@
We explore the effect of history length $m$ on learning the inverse observation operator and its impact on state estimation accuracy. According to the theory in (Liu et al., 2022), the history length can be chosen as $m\propto\log(n_{s})$ [^3] Please note that the result omits the class of systems with exponential dependency on the history length.. The ablation study is conducted by scaling $m$ proportionally to $m\approx C\log{n_{s}}$ where the constant $C$ was adjusted. The dimensions $d_{s},d_{o},d_{h}$ are fixed to be the same as the experiments in Subsection 4.1. Table 3 (right) shows that incorporating history ($C>0$) significantly improves the accuracy of the state estimation, with the NRMSE decreasing as $C$ increases. However, improvements become marginal when increasing $C$ beyond a certain point (around $C=4$). This indicates a trade-off, where increasing the history length beyond a certain threshold yields little additional benefit in state estimation.
@@
This experiment compares DFs with varying dimensions $d_{s}=20,40,60,80$ to a Gaussian kernel feature. To scale up the Gaussian kernel, we apply Nyström approximation and kernel PCA (See Appendix E.5 for details). All three features are Gaussian kernels. The preimage of the system state is learned using kernel ridge regression on the low-dimensional representations. As shown in Table 3 (left), Gaussian kernel performance is close to the best DFs in $n_{s}=40$ but degrades in $n_{s}=80$. In a lower-dimensional problem, the Gaussian kernel features are more robust than DFs, but their performance can degrade with increasing system dimensionality and complexity. DFs with $d_{s}=60$ and $120$ perform consistently well, whereas $d_{s}=20$ and $40$ underperform, indicating trade-offs between data size, model complexity, and system dimension.
Table 3: Comparison of different features and history lengths with NRMSE (%) as the metric for state estimation accuracy. GK denotes the Gaussian kernel.
Feature Dimension (NRMSE %) History Length (NRMSE %)
$d_{s}$ $n_{s}=40$ $n_{s}=80$ $C$ $n_{s}=40$ $n_{s}=80$
$20$ $16.7\pm 2.1$ $17.3\pm 2.6$ 00 $11.7\pm 3.4$ $10.8\pm 2.6$
$40$ $14.1\pm 1.3$ $16.8\pm 2.2$ $1$ $9.3\pm 2.8$ $8.2\pm 1.7$
$60$ 8.3 $\pm$ 0.9 $11.4\pm 0.9$ $2$ $7.7\pm 1.4$ $7.4\pm 0.9$
$120$ $9.7\pm 0.7$ 10.9 $\pm$ 0.9 $4$ $7.7\pm 0.9$ $7.1\pm 0.9$
GK $8.4\pm$ 0.5 $11.7\pm 1.4$ $8$ $\mathbf{7.5}\pm\mathbf{0.5}$ $\mathbf{7.0}\pm\mathbf{0.8}$
@@ 5 Conclusion
In this paper, we propose Tensor-Var, a framework that uses kernel conditional mean embedding to linearize nonlinear models in 4D-Var, making optimization tractable and efficient. By learning adaptive deep features, Tensor-Var addresses the scalability typically associated with traditional kernel methods. Our inverse observation operator, which incorporates historical observations, improves accuracy and robustness with incomplete observations. Experiments on two chaotic systems and global weather forecasting show that Tensor-Var outperforms state-of-the-art hybrid ML-DA models in both accuracy and efficiency.
Limitations and future work.
@@
Tensor-Var requires access to system states to learn dynamics and observation models, which may not be feasible in practice. Future work will focus on learning these models directly from observations and calibrating dynamics in feature space. Additionally, the simplified error covariance used in 4D-Var may underestimate system correlations; refining these within Tensor-Var could improve assimilation and performance.
Acknowledgments
The authors thank the reviewers and the ICML 2025 program committee for their insightful feedback and constructive suggestions that have helped improve this work. We also acknowledge the support from the University College London (Dean’s Prize, Chadwick Scholarship), the Engineering and Physical Sciences Research Council projects (EP/W007762/1), the United Kingdom Atomic Energy Authority (NEPTUNE 2057701-TN-03), and the Royal Academy of Engineering (IF-2425-19-AI165).
Impact Statement
This paper aims to advance machine learning applications in studying the long-term properties of large-scale chaotic systems. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Table of Notations
Notations Meaning
$S,O,H$ random variables of state, observation, and history.
$s,o,h$ samples of state, observation, and history.
$\mathcal{S}$ domain of state space.
$\mathcal{O}$ domain of observation space.
$n$ dimension of original spaces with specified subscripts.
$\mathbb{P}$ probability distribution with specified subscripts.
$\phi_{S}(s)=k_{S}(s,\cdot)$ kernel feature map for state.
$\phi_{O}(o)=k_{O}(o,\cdot)$ kernel feature map for observation.
$\phi_{H}(h)=k_{H}(h,\cdot)$ kernel feature map for history.
$\mathbb{H}_{S},\mathbb{H}_{O},\mathbb{H}_{H}$ induced kernel feature space for $S,O$ and $H$.
$z$ indicate the elements in feature space.
$\mathcal{C}$ Conditional mean embedding (CME) operators with specified subscripts.
$B,R,Q$ error matrices for background, observation, and dynamical model in original space.
$\mathcal{B},\mathcal{R},\mathcal{Q}$ error matrices for background, observation, and dynamical model in feature space.
$\phi_{\theta_{S}}$ deep feature for state with neural network parameters $\theta_{S}$.
$\phi^{\dagger}_{\theta^{\prime}_{S}}$ inverse feature for state with neural network parameters $\theta_{S}$.
$\phi_{\theta_{O}}$ deep feature for state with neural network parameters $\theta_{O}$.
$\phi_{\theta_{H}}$ deep feature for state with neural network parameters $\theta_{H}$.
$d$ dimension of feature spaces with specified subscripts.
Appendix B Variational Data Assimilation Optimization
A sequence $\{x_{k}\}$ generated by an optimization algorithm converges linearly to $x^{*}$ if there exists $C\in(0,1)$ such that $\|x_{k+1}-x^{*}\|\leq C\|x_{k}-x^{*}\|$, implying $\|x_{k}-x^{*}\|=\mathcal{O}(C^{k})$. Convergence is sublinear if $\|x_{k}-x^{*}\|\to 0$ as $k\to\infty$ but slower than linear, often expressed as $\|x_{k}-x^{*}\|=\mathcal{O}(1/k^{\alpha})$ for $\alpha>0$. Superlinear convergence occurs when $\lim_{k\to\infty}\|x_{k+1}-x^{*}\|/\|x_{k}-x^{*}\|=0$, with quadratic convergence as a special case where $\|x_{k+1}-x^{*}\|\leq C\|x_{k}-x^{*}\|^{2}$.
B.1 Analysis of Optimization
Tensor-Var.
Our approach based on CME is theoretically linear in the learned feature space. Under this condition, it formulates a convex quadratic programming (QP) problem with linear dynamic constraints, as illustrated in Equation (5). From the convergence rates of the Newton optimization method (Wright, 2006), the algorithm exhibits a linear convergence rate in worst-case and a superlinear convergence rate in best-case. Consequently, the number of iterations required to achieve an $\epsilon$-error ranges between $\mathcal{O}(\log(1/\epsilon))$ and $\mathcal{O}(\log\log(1/\epsilon))$. In terms of computational complexity, the QP problem features an invariant Hessian matrix. This invariance allows us to compute the Hessian matrix only once. Incorporating the iteration complexity, the total computational complexity of Tensor-Var is efficiently bounded between $\mathcal{O}\left(Td_{s}\log(1/\epsilon)\right)$ and $\mathcal{O}\left(Td_{s}\log\log(1/\epsilon)\right)$. This demonstrates that Tensor-Var achieves efficient computation by minimizing redundant Hessian calculations and leveraging the convergence properties of the Newton method. It should be noted that the optimization error in Tensor-Var is in the feature space, not the state space. The true error is quantified as $(\epsilon+\hat{\text{err}})$ it decoding back, where $\hat{\text{err}}$ represents the reconstruction error.
Latent 4D-Var.
Unlike our linearized CME model, the latent 4D-Var approach (Cheng et al., 2024) directly constructs a neural network-based surrogate model using an autoencoder, thereby no longer depending on computationally intensive numerical models. In this framework, the inherent nonlinearity and non-convexity of the optimization problem remain unchanged. According to optimization theory, the L-BFGS algorithm achieves a worst-case sublinear convergence rate and a best-case superlinear convergence rate (Boyd & Vandenberghe, 2004; Wright, 2006). The best computational complexity can only be achieved only if the problem exhibits strong local convexity around the initial point. Consequently, the number of iterations required to achieve an $\epsilon$-error ranges between $\mathcal{O}(1/\epsilon^{2})$ and $\mathcal{O}(\log\log(1/\epsilon))$. Regarding computational complexity, latent 4D-Var does not decompose the quadratic programming problem into $T$ separate subproblems. Instead, it concatenates the $T$ features into a high-dimensional vector of size $Td_{s}$. At each time step, the Hessian matrix is updated using the BFGS algorithm. Considering the complexity of the iteration, the total computational complexity ranges from $\mathcal{O}((Td_{s})^{2}/\epsilon^{2})$ to $\mathcal{O}((Td_{s})^{2}\log\log(1/\epsilon))$ which makes it computationally expensive. Since optimization is performed in the latent feature space, the true error is quantified as $(\epsilon+\hat{\text{err}})$ upon decoding, where $\hat{\text{err}}$ represents the reconstruction error.
4D-Var.
The original 4D-Var, when solved using the Newton method, exhibits a similar convergence rate, ranging from sublinear to superlinear. However, the number of iterations scales with the state dimension $n_{s}$ rather than the latent space dimension $d_{s}$, leading to a total complexity of $\mathcal{O}((Tn_{s})^{2}/\epsilon^{2})$ or $\mathcal{O}\big{(}(Tn_{s})^{2}\log\log(1/\epsilon)\big{)}$ to achieve an $\epsilon$-error. When the adjoint model is available, the total computational complexities can be further reduced to $\mathcal{O}(Tn_{s}/\epsilon^{2})$ or $\mathcal{O}\big{(}Tn_{s}\log\log(1/\epsilon)\big{)}$ through forward and adjoint computations, each requiring $\mathcal{O}(n_{s})$ (Givoli, 2021; Wright, 2006). Since the 4D-Var is directly solved on the original space, there is no reconstruction error that needs to be counted.
Frerix et al., 2021.
Frerix et al. (2021) propose learning an approximate inverse observation operator using deep learning and transform this problem into the physical state space. The 4D-Var problem is then solved using a standard Newton solver, resulting in convergence properties and computational complexities similar to those of the standard 4D-Var without an adjoint model such as $\mathcal{O}((Tn_{s})^{2}/\epsilon^{2})$ or $\mathcal{O}\big{(}(Tn_{s})^{2}\log\log(1/\epsilon)\big{)}$. In practice, (Frerix et al., 2021) argues that solving 4D-Var in a physical state space with an inverse observation operator leads to better initialization, as inverse mapping provides an improved starting point for optimization. This improved initialization results in a better-conditioned optimization landscape, i.e., strongly local convex and achieves a superlinear convergence rate. Our numerical experiments further confirm that their method achieves faster convergence with fewer iterations compared to the standard 4D-Var model (with or without an adjoint model), as shown in Table 2.
A comprehensive summary of these analyses and findings is provided in Table 4.
Table 4: Full comparison of Tensor-Var with other algorithms. Here, $\epsilon$ is the error threshold and $T$ is the length of the assimilation window. The best computational complexity can only be achieved only if the problem exhibits strong local convexity around the initial point.
Space Transformation Non-Space Transformation
Methods Tensor-Var Latent 4D-Var 4D-Var (w/o or w/ Adjoint) Frerix et al. (2021)
Model-based
Convexity
Convergence Rate linear to superlinear sublinear to linear sublinear to linear sublinear to linear
Iteration Time $\mathcal{O}\left(\log\left(1/\epsilon\right)\right)$ $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}\left(1/\epsilon^{2}\right)$ $\mathcal{O}\left(1/\epsilon^{2}\right)$
Worst Computational
Complexity
$\mathcal{O}\left(Td_{s}\log\left(1/\epsilon\right)\right)$ $\mathcal{O}\left((Td_{s})^{2}/\epsilon^{2}\right)$ w/o Adjoint: $\mathcal{O}\left((Tn_{s})^{2}/\epsilon^{2}\right)$
w/ Adjoint: $\mathcal{O}\left(Tn_{s}/\epsilon^{2}\right)$
$\mathcal{O}\left((Tn_{s})^{2}/\epsilon^{2}\right)$
Best Computational
Complexity
$\mathcal{O}\left(Td_{s}\log\log\left(1/\epsilon\right)\right)$ $\mathcal{O}\left((Td_{s})^{2}\log\log(1/\epsilon)\right)$ w/o Adjoint: $\mathcal{O}\left((Tn_{s})^{2}\log\log\left(1/\epsilon\right)\right)$
w/ Adjoint: $\mathcal{O}\left(Tn_{s}\log\log\left(1/\epsilon\right)\right)$
$\mathcal{O}\left((Tn_{s})^{2}\log\log\left(1/\epsilon\right)\right)$
Appendix C Theoretical Analysis